# Leveraging (Biased) Information: Multi-armed Bandits with Offline Data

### v1

@@ B.3 Proof for instance independent regret upper bounds
$\leq K\Delta_{\text{max}}+\sum^{T}_{t=K+1}\left[\min\{\mu^{(\text{on})}(A_{t})+2\text{rad}_{t}(A_{t}),\mu^{(\text{on})}(A_{t})+2\text{rad}^{\text{S}}_{t}(A(t))\}-\mu^{(\text{on})}(A_{t})\right]$ (29)
$\leq K\Delta_{\text{max}}+2\min\left\{\sum^{T}_{t=K+1}\text{rad}_{t}(A_{t}),\sum^{T}_{t=K+1}\text{rad}^{\text{S}}_{t}(A_{t})\right\}.$ (30)
Step (29) is by applying Lemma 4.1, and noting that $\mu^{(\text{off})}(a)-\mu^{\text{(on)}}(a)\leq V(a)$ in the upper bound of $\text{UCB}^{\text{S}}_{t}(a)$ in the Lemma.
The conventional analysis shows that $\sum^{T}_{t=K+1}\text{rad}_{t}(A_{t})=O(\sqrt{KT\log(T/\delta)})$. We next focus on upper bounding
@@
We show (31) by establishing two major claims:
Claim 1: Suppose that $(N^{*}_{T}(a))_{a\in{\cal A}}\in\mathbb{N}_{\geq 0}^{\cal A}$ is an optimal solution to the following maximization problem
[…]
s.t. $\sum_{a\in{\cal A}}N_{T}(a)=T$ ,
[…]
Then it must be the case that $N^{*}_{T}(a)\leq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}$ for all $a\in{\cal A}$, where we recall that $\tau_{*}$ is the optimum of (LP).
Claim 2: . For an optimal solution $(\tau_{*},\{n_{*}(a)\}_{a\in{\cal A}})$ to (LP), it holds that
$n_{*}(a)=\max\{\tau_{*}-T_{\text{S}}(a),0\}\text{ for every arm $a$.}$ (32)
We see that Claim 1 is immediately useful for proving (31), since then
@@
where the first inequality is by the optimality of $(N^{*}_{T}(a))_{a\in{\cal A}}$ to (IP), and the fact that any realized $(N_{T}(a))_{a\in{\cal A}}$ must be feasible to (IP). The second inequality is a direct consequence of Claim 1. Claim 2 is an axuiliary claim for proving Claim 1. We postpone the proofs of these two claims to the end of the proof of the Theorem.
Given (31), for each $a\in{\cal A}$ we then have
@@
hence $\sum_{a\in{\cal A}}\sum^{N_{T}(a)}_{n=1}\sqrt{\frac{8\log(T/\delta)}{n+T_{\text{S}}(a)}}=O(T\sqrt{\log(T/\delta)/\tau_{*}})$. Altogether, the Theorem is proved after we provide the proofs to Claims 1, 2:
Proving Claim 1 We establish the claim by a contradiction argument. Suppose there exists an arm $a$ such that $N^{*}_{T}(a)\geq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}+1$. Firstly, we assert that there must exist another arm $a^{\prime}\in{\cal A}\setminus\{a\}$ such that $N^{*}_{T}(a^{\prime})\leq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a^{\prime}),0\}-1$. If not, we then have $N^{*}_{T}(a)\geq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}+1$ and $N^{*}_{T}(a^{\prime\prime})\geq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a^{\prime\prime}),0\}$ for all $a^{\prime\prime}\in{\cal A}$, but then
$\sum_{k\in{\cal A}}N^{*}_{T}(k)>\sum_{k\in{\cal A}}\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(k),0\}\geq\sum_{k\in{\cal A}}\max\{\tau_{*}-T_{\text{S}}(k),0\}\overset{(\dagger)}{=}\sum_{k\in{\cal A}}n_{*}(k)=T,$
violating the constraint $\sum_{k\in{\cal A}}N^{*}_{T}(k)=T$. Note that the equality ()($\dagger$)( † ) is by Claim 2. Thus, the claimed arm $a^{\prime}$ exists. To this end, the condition $N^{*}_{T}(a^{\prime})\leq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a^{\prime}),0\}-1$ implies that $\lceil\tau_{*}\rceil>T_{\text{S}}(a^{\prime})$, since otherwise $N^{*}_{T}(a^{\prime})\leq-1$, which violates the non-negativity constraint on $N^{*}_{T}(a)$. Altogether we have established the existence of two distinct arms $a,a^{\prime}\in{\cal A}$ such that
$N^{*}_{T}(a)+T_{\text{S}}(a)$ $\geq\max\{\lceil\tau_{*}\rceil,T_{\text{S}}(a)\}+1\text{, in particular }N^{*}_{T}(a)\geq 1,$
$N^{*}_{T}(a^{\prime})+T_{\text{S}}(a^{\prime})$ $\leq\max\{\lceil\tau_{*}\rceil,T_{\text{S}}(a^{\prime})\}-1=\lceil\tau_{*}\rceil-1.$
To establish the contradiction argument, consider another solution $(\tilde{N}_{T}(k))_{k\in{\cal A}}$ to the displayed optimization problem in the claim, where
$\tilde{N}_{T}(k)=\begin{cases}N^{*}_{T}(k)-1&\quad\text{if }k=a\\ N^{*}_{T}(k)+1&\quad\text{if }k=a^{\prime}\\ N^{*}_{T}(k)&\quad\text{if }k\in{\cal A}\setminus\{a,a^{\prime}\}.\end{cases}$
By the property that $N^{*}_{T}(a)\geq 1$, $\tilde{N}_{T}(a)\geq 0$, and $(\tilde{N}_{T}(k))_{k\in{\cal A}}$ is a feasible solution. But then we have
$\sum_{a\in{\cal A}}\sum^{\tilde{N}_{T}(a)}_{n(a)=1}\sqrt{\frac{1}{n(a)+T_{\text{S}}(a)}}-\sum_{a\in{\cal A}}\sum^{N^{*}_{T}(a)}_{n(a)=1}\sqrt{\frac{1}{n(a)+T_{\text{S}}(a)}}$
$=$ $\sqrt{\frac{1}{N^{*}_{T}(a^{\prime})+T_{\text{S}}(a^{\prime})+1}}-\sqrt{\frac{1}{N^{*}_{T}(a)+T_{\text{S}}(a)}}$
$\geq$ $\sqrt{\frac{1}{\lceil\tau_{*}\rceil}}-\sqrt{\frac{1}{\max\{\lceil\tau_{*}\rceil,T_{\text{S}}(a)\}+1}}>0,$ (34)
which contradicts the assumed optimality of $(N^{*}_{T}(a))_{a\in{\cal A}}$. Thus Claim 1 is shown.
Proving Claim 2 . Now, since $(\tau_{*},\{n_{*}(a)\}_{a\in{\cal A}})$ is feasible to the (LP), the constraints in the (LP) give us that $n_{*}(a)\geq\max\{\tau_{*}-T_{\text{S}}(a),0\}$. We establish our asserted equality by a contradition argument. Suppose there is an arm $a^{\prime}$ such that $n_{*}(a^{\prime})-\max\{\tau_{*}-T_{\text{S}}(a^{\prime}),0\}=\epsilon>0$. Then it can be verified that the solution $(\tau^{\prime},\{n^{\prime}(a)\}_{a\in{\cal A}})$ defined as $\tau^{\prime}=\tau_{*}+\epsilon/K$, $n^{\prime}(a)=n_{*}(a)+\epsilon/K$ for all $a\in{\cal A}\setminus\{a^{\prime}\}$ and $n^{\prime}(a^{\prime})=n_{*}(a)-\frac{K-1}{K}\epsilon$ is also a feasible solution to (LP). But then $\tau^{\prime}>\tau_{*}$, which contradicts with the optimality of $(\tau_{*},\{n_{*}(a)\}_{a\in{\cal A}})$. Thus (32) is shown.
Appendix C Proofs for Regret Lower Bounds
C.1 Notational Set Up and Auxiliary Results
@@
Notational Set Up. Consider a non-anticipatory policy $\pi$ and an instance $I$ with reward distribution $P$. We denote $\rho_{P,\pi}$ as the joint probability distribution function on $S,A_{1},R_{1},\ldots,A_{T},R_{T}$, the concatenation of the offline dataset $S$ and the online trajectory under policy $\pi$ on instance $I$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable event $E$, we denote $\Pr_{P,\pi}(E)$ as the probability of $E$ holds under $\rho_{P,\pi}$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable random variable $Y$, we denote $\mathbb{E}_{P,\pi}[Y]$ as the expectation of $Y$ under the joint probability distribution function $\rho_{P,\pi}$. In our analysis, we make use of $Y$ being $\text{Reg}_{T}(\pi,P)$ or $N_{T}(a)$ for an arm $a$. To lighten our notational burden, we abbreviate $\mathbb{E}_{P,\pi}[\text{Reg}_{T}(\pi,P)]$ as $\mathbb{E}[\text{Reg}_{T}(\pi,P)]$.
@@
Auxiliary Results To establish the regret lower bounds and the impossibility result, we need three auxiliary results, namely Claim 1 and Theorems C.1, C.2. Claim 1 is on the KL-divergence between Gaussian random variables:
Claim 1.
@@
The following, dubbed Bretagnolle–Huber inequality, is extracted from Theorem 14.2 from (Lattimore & Szepesvári, 2020):
Theorem C.1.
@@
Lastly, the derivation of the chain rule in Theorem C.2 largely follows from the derivation of Lemma 15.1 in (Lattimore & Szepesvári, 2020):
Theorem C.2.
@@
We provide a proof to theorem C.2 for completeness sake:
Proof of Theorem C.2.
The proof largely follow the well-known chain rule in the multi-armed bandit literature, for example see Lemma 15.1 in (Lattimore & Szepesvári, 2020). We start by explicitly expressing the joint probability function $\rho_{P,\pi}$ on
@@
$\quad\log\left(\frac{P^{\text{(on)}}_{a_{t}}(r_{t})}{Q^{\text{(on)}}_{a_{t}}(r_{t})}\right)\text{ d}x\text{ d}r_{1}\ldots\text{ d}r_{t}$
@@
×$\quad\times\pi_{t}(a_{t}|x,a_{1},r_{1},\ldots,a_{t-1},r_{t-1})\text{ d}x\text{ d}r_{1}\ldots\text{ d}r_{t-1}\Big{\}}\cdot\int_{r_{t}}P^{\text{(on)}}_{a_{t}}(r_{t})\log\left(\frac{P^{\text{(on)}}_{a_{t}}(r_{t})}{Q^{\text{(on)}}_{a_{t}}(r_{t})}\right)\text{ d}r_{t}$
@@
C.2 Proof for Proposition 3.1, the Impossibility Result
By the condition $C<T^{\epsilon}/(4\log T)$, we know that $\frac{1}{\sqrt{C\log T}T^{\beta-(\epsilon/2)}}-\frac{1}{T^{\beta}}>\frac{1}{T^{\beta}}>0$, so in the instance with $Q$, arm 2 is the unqiue optimal arm. Consider the event $E=\{N_{T}(2)>T/2\}$. Now,
$\mathbb{E}[\text{Reg}(\pi,P)]+\mathbb{E}[\text{Reg}(\pi,Q)]$
$\geq$ $\frac{1}{T^{\beta}}\cdot\frac{T}{2}\cdot\Pr_{P,\pi}(E)+\left[\frac{1}{\sqrt{C\log T}T^{\beta-(\epsilon/2)}}-\frac{1}{T^{\beta}}\right]\cdot\frac{T}{2}\Pr_{Q,\pi}(E^{C})$
$\geq$ $\frac{T^{1-\beta}}{4}\cdot\exp\left[-\mathbb{E}_{\pi,P}[N_{T}(2)]\cdot\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})-T_{S}(2)\cdot\text{KL}(P^{(\text{off})}_{2},Q^{(\text{off})}_{2})\right]$ (38)
$=$ $\frac{T^{1-\beta}}{4}\cdot\exp\left[-\mathbb{E}_{\pi,P}[N_{T}(2)]\cdot\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})\right].$ (39)
Step (38) is again by the chain rule in Theorem C.2. Step (39) is because $P^{\text{(off)}}=Q^{\text{(off)}}$. It is evident that
$\text{KL}(P^{(\text{on})}_{2},Q^{(\text{on})}_{2})=\frac{1}{2CT^{2\beta-\epsilon}\log T},$
and by the Claim assumption that $\text{Reg}(\pi,P)=(1/T^{\beta})\mathbb{E}_{\pi,P}[N_{T}(2)]\leq CT^{\beta-\epsilon}\log T$, we have
$\mathbb{E}_{\pi,P}[N_{T}(2)]\leq CT^{2\beta-\epsilon}\log T,$
which leads to
Again by the Proposition assumption $\mathbb{E}[\text{Reg}(\pi,P)]\leq CT^{\beta-\epsilon}\log T$, we arrive at the claimed inequality, and the Proposition is proved.
C.3 Proof for instance dependent regret lower bound
C.3.1 Proof for Theorem 4.7

### v2

@@ C.2.1  Proof for Theorem 4
$\leq K\Delta_{\text{max}}+\sum^{T}_{t=K+1}\left[\min\{\mu^{(\text{on})}(A_{t})+2\text{rad}_{t}(A_{t}),\mu^{(\text{on})}(A_{t})+2\text{rad}^{\text{S}}_{t}(A(t))\}-\mu^{(\text{on})}(A_{t})\right]$ (40a)
$\leq K\Delta_{\text{max}}+2\min\left\{\sum^{T}_{t=K+1}\text{rad}_{t}(A_{t}),\sum^{T}_{t=K+1}\text{rad}^{\text{S}}_{t}(A_{t})\right\}.$ (40b)
(40a) is by applying Lemma 1, and $\mu^{(\text{off})}(a)-\mu^{\text{(on)}}(a)\leq V(a)$ in the upper bound of $\text{UCB}^{\text{S}}_{t}(a)$. The conventional analysis shows that $\sum^{T}_{t=K+1}\text{rad}_{t}(A_{t})=O(\sqrt{KT\log(T/\delta)})$. We focus on:
@@
Consider the following integer programming problem, and denote the optimal solution as $(N^{*}_{T}(a))_{a\in{\cal A}}$ where $(N^{*}_{T}(a))_{a\in{\cal A}}\in\mathbb{N}_{\geq 0}^{\cal A}$:
$\text{(IP)}:\max_{(N_{T}(a))_{a\in{\cal A}}}$ $\sum_{a\in{\cal A}}\sum^{N_{T}(a)}_{n(a)=1}\sqrt{\frac{1}{n(a)+T_{\text{S}}(a)}}$
s.t. $\sum_{a\in{\cal A}}N_{T}(a)=T$
$N_{T}(a)\in\mathbb{N}_{\geq 0},\ \forall a\in\mathcal{A}.$
Then Lemma 5 (with $\rho=1$) is immediately useful for proving (41), since
@@
where the first inequality is by the optimality of $(N^{*}_{T}(a))_{a\in{\cal A}}$ to (IP), and the fact that any realized $(N_{T}(a))_{a\in{\cal A}}$ must be feasible to (IP). The second inequality is a direct consequence of Lemma 5. For each $a\in{\cal A}$ we then have
@@
hence $\sum_{a\in{\cal A}}\sum^{N_{T}(a)}_{n=1}\sqrt{\frac{8\log(T/\delta)}{n+T_{\text{S}}(a)}}=O(T\sqrt{\log(T/\delta)/\tau_{*}})$. Altogether, the Theorem is proved.
C.3  Proof for Instance-independent bound for Combinatorial Bandit
C.3.1  Proof for Theorem 6
Denote $\text{minUCB}_{t}=\left(\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}\right)_{a\in\mathcal{A}}$. Conditioned on $\cap^{T}_{t=1}\xi_{t}$, we have
$\text{Reg}_{T}\$$=\sum_{t=1}^{T}\left(\alpha\cdot\beta\cdot\text{r}^{*}_{\boldsymbol{\mu}^{\text{(on)}}}-\mathbb{E}\left[\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A_{t})\right]\right)$
$=\sum_{t=1}^{T}\left(\alpha\cdot\beta\cdot\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A^{*}_{\boldsymbol{\mu}^{\text{(on)}}})-\mathbb{E}\left[\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A_{t})\right]\right)$ (43a)
$\leq\sum_{t=1}^{T}\left(\alpha\cdot\beta\cdot\text{r}_{\text{minUCB}_{t}}(A^{*}_{\boldsymbol{\mu}^{\text{(on)}}})-\mathbb{E}\left[\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A_{t})\right]\right)$ (43b)
$\leq\sum_{t=1}^{T}\left(\alpha\cdot\beta\cdot\text{r}_{\text{minUCB}_{t}}^{*}-\mathbb{E}\left[\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A_{t})\right]\right)$
$\leq K\Delta_{\max}+\sum_{t=1}^{T}\left(\mathbb{E}\left[\text{r}_{\text{minUCB}_{t}}(A_{t})\right]-\mathbb{E}\left[\text{r}_{\boldsymbol{\mu}^{\text{(on)}}}(A_{t})\right]\right)$ (43c)
$\leq K\Delta_{\max}+\gamma\cdot\sum_{t=1}^{T}\left(\max_{a\in A_{t}}\left\{\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}-\boldsymbol{\mu}^{\text{(on)}}(a)\right\}\right)^{\rho}$ (43d)
$=K\Delta_{\max}+\gamma\cdot\sum_{t=1}^{T}\min\left\{\text{UCB}_{t}(a_{t})-\boldsymbol{\mu}^{\text{(on)}}(a_{t}),\text{UCB}^{\text{S}}_{t}(a)-\boldsymbol{\mu}^{\text{(on)}}(a_{t})\right\}^{\rho}$ (43e)
$\leq K\Delta_{\max}+2^{\rho}\gamma\cdot\sum_{t=1}^{T}\min\left\{\text{rad}_{t}(a_{t}),\text{rad}^{\text{S}}_{t}(a_{t})\right\}^{\rho}$ (43f)
$=K\Delta_{\max}+2^{\rho}\gamma\cdot\sum_{t=1}^{T}\min\left\{\text{rad}_{t}(a_{t})^{\rho},\text{rad}^{\text{S}}_{t}(a_{t})^{\rho}\right\}$
$\leq K\Delta_{\max}+2^{\rho}\gamma\cdot\min\left\{\sum_{t=1}^{T}\text{rad}_{t}(a_{t})^{\rho},\sum_{t=1}^{T}\text{rad}^{\text{S}}_{t}(a_{t})^{\rho}\right\}.$
Here, (43a) comes from the definition of $\text{r}_{\boldsymbol{u}}^{*}$ and $A^{*}_{\boldsymbol{u}}$. (43b) comes from the monotonicity of $\text{r}_{\boldsymbol{u}}(\cdot)$ and Lemma 1. (43c) comes from the selection rule (19) for $A_{t}$. By the definition of $(\alpha,\beta)$ optimization oracle, we have $\mathbb{E}[\text{r}_{\text{minUCB}_{t}}(A_{t})]\geq\alpha\cdot\beta\cdot\text{r}_{\text{minUCB}_{t}}^{*}$. (43d) comes from the bounded smoothness of $\text{r}_{\boldsymbol{u}}(\cdot)$ and the expression of $f(\cdot)$. In (43e) $a_{t}=\mathop{\arg\max}_{a\in A_{t}}\left\{\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}-\boldsymbol{\mu}^{\text{(on)}}(a)\right\}$. (43f) is by Lemma 1. Denote
$N_{t}^{\max}(a)=\max\left\{1,\sum_{s=1}^{t-1}\boldsymbol{1}\left\{a\in\mathop{\arg\max}_{a^{\prime}\in A_{t}}\left\{\min\left\{\text{UCB}_{t}(a^{\prime}),\text{UCB}^{\text{S}}_{t}(a^{\prime})\right\}-\boldsymbol{\mu}^{\text{(on)}}(a^{\prime})\right\}\right\}\right\}$
be the number of times that $a$ becomes the maximum gap arm. Now we analyze the two terms in $\min$, respectively. Firstly, by the expression of $\text{rad}_{t}(a)$,
$\sum_{t=1}^{T}\text{rad}_{t}(a_{t})^{\rho}$ $\leq\sqrt{2\log(4KT^{3}/\delta)}\sum_{t=1}^{T}N_{t}(a_{t})^{-\rho/2}$
$\leq(2\log(4KT^{3}/\delta))^{\rho/2}\sum_{t=1}^{T}N_{t}^{\max}(a_{t})^{-\rho/2}$ (44a)
$\leq(2\log(4KT^{3}/\delta))^{\rho/2}\sum_{a\in\mathcal{A}}\sum_{n=1}^{N_{T}^{\max}(a)}n^{-\rho/2}$ (44b)
$\leq\frac{1}{1-\rho/2}(2\log(4KT^{3}/\delta))^{\rho/2}\sum_{a\in\mathcal{A}}N_{T}^{\max}(a)^{1-\rho/2}$
$\leq\frac{1}{1-\rho/2}(2\log(4KT^{3}/\delta))^{\rho/2}\cdot K^{\rho/2}\cdot T^{1-\rho/2},$ (44c)
where (44a) comes from the fact that $a_{t}\in A_{t}$, hence $N_{t}(a)\geq N_{t}^{\max}(a)$. (44b) comes from rearanging the sum. (44c) comes from Jensen Inequality and the fact that $\sum_{a\in\mathcal{A}}N_{T}^{\max}(a)=T$. Secondly, by the expression of $\text{rad}^{\text{S}}_{t}(a)$,
$\sum^{T}_{t=1}\text{rad}^{\text{S}}_{t}(a_{t})^{\rho}$ $=\sum^{T}_{t=1}\left[\sqrt{\frac{2\log(2t/\delta_{t})}{N_{t}(a_{t})+T_{\text{S}}(a_{t})}}+\frac{T_{\text{S}}(a_{t})}{N_{t}(a_{t})+T_{\text{S}}(a_{t})}\cdot V(a_{t})\right]^{\rho}$
$\leq V_{\max}^{\rho}\cdot T+(2\log(4KT^{3}/\delta))^{\rho/2}\sum_{a\in{\cal A}}\sum^{N_{T}(a)}_{n=1}(n+T_{\text{S}}(a))^{-\rho/2}.$
Similarly, suppose $(N_{T}^{*}(a))_{a\in\mathcal{A}}$ be an optimal solution for IP (38), then we claim that
$\sum_{a\in{\cal A}}\sum^{N_{T}(a)}_{n(a)=1}(n(a)+T_{\text{S}}(a))^{-\rho/2}\leq\sum_{a\in{\cal A}}\sum^{N^{*}_{T}(a)}_{n(a)=1}(n(a)+T_{\text{S}}(a))^{-\rho/2}\leq\sum_{a\in{\cal A}}\sum^{\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}}_{n(a)=1}(n(a)+T_{\text{S}}(a))^{-\rho/2}.$
For each $a\in{\cal A}$ we then have
$\sum^{\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}}_{n=1}(n+T_{\text{S}}(a))^{-\rho/2}\leq\frac{\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}}{\lceil\tau_{*}\rceil}\sum^{\lceil\tau_{*}\rceil}_{t=1}t^{-\rho/2}\leq\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}\cdot\frac{1}{1-\rho/2}\cdot\tau_{*}^{-\rho/2}.$
Combining above, we have
$\sum_{a\in{\cal A}}\sum^{\max\{\lceil\tau_{*}\rceil-T_{\text{S}}(a),0\}}_{n=1}(n+T_{\text{S}}(a))^{-\rho/2}\leq\sum_{a\in{\cal A}}(n_{*}(a)+1)\cdot\frac{1}{1-\rho/2}\cdot\tau_{*}^{-\rho/2}\leq T\cdot\frac{2}{1-\rho/2}\cdot\tau_{*}^{-\rho/2}.$
Hence
$\sum^{T}_{t=1}\text{rad}^{\text{S}}_{t}(a_{t})^{\rho}\leq O\left(\left(V_{\max}^{\rho}+(\log(T/\delta))^{\rho/2}\cdot\tau_{*}^{-\rho/2}\right)\cdot T\right).$
Altogether, the Theorem is proved.
C.4  Proof for instance-independent bound for Combinatorial Bandits with Linear Rewards
C.4.1  Proof for Theorem 22
Similarly, conditioned on $\cap^{T}_{t=1}\xi_{t}$, we have
$\text{Reg}_{T}$ $=\sum_{t=1}^{T}\left(\sum_{a\in A_{*}}\mu^{\text{(on)}}(a)-\sum_{a\in A_{t}}\mu^{\text{(on)}}(a)\right)$ (46a)
$\leq K\Delta_{\max}+\sum_{t=1}^{T}\sum_{a\in A_{t}}\left(\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}-\mu^{\text{(on)}}(a)\right)$ (46b)
$\leq K\Delta_{\max}+\min\left\{\sum_{t=1}^{T}\sum_{a\in A_{t}}\left(\text{UCB}_{t}(a)-\mu^{\text{(on)}}(a)\right),\sum_{t=1}^{T}\sum_{a\in A_{t}}\left(\text{UCB}^{\text{S}}_{t}(a)-\mu^{\text{(on)}}(a)\right)\right\}$
$\leq K\Delta_{\max}+2\min\left\{\sum_{t=1}^{T}\sum_{a\in A_{t}}\text{rad}_{t}(a),\sum_{t=1}^{T}\sum_{a\in A_{t}}\text{rad}_{t}^{\text{S}}(a_{t})\right\}.$
where (46a) comes from the definition of $\text{r}_{\boldsymbol{u}}(A)=\sum_{a\in A}u(a)$. (46b) comes from the following that
$\sum_{a\in A_{*}}\mu^{\text{(on)}}(a)\leq\sum_{a\in A_{*}}\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}\leq\sum_{a\in A_{t}}\min\left\{\text{UCB}_{t}(a),\text{UCB}^{\text{S}}_{t}(a)\right\}.$
By conventional analysis, we can obtain
$\sum_{t=1}^{T}\sum_{a\in A_{t}}\text{rad}_{t}(a)\leq O\left(\sqrt{mKT\log(T/\delta)}\right).$
By the analysis similar as the proof of Theorem 4, we can obtain
$\sum_{t=1}^{T}\sum_{a\in A_{t}}\text{rad}_{t}^{\text{S}}(a_{t})\leq O\left(\left(\sqrt{\frac{\log(T/\delta)}{\tau_{*}^{\text{C}}}}+V_{\text{max}}\right)\cdot mT\right).$
Altogether, the Theorem is proved.
C.4.2  Discussions: (21) v.s. (20)
In the case of combinatorial bandits with linear rewards, if we directly apply Theorem 6, we can obtain bound (20). However, in this special case, a tighter bound can be achieved. The key observation comes from the structure of the reward function. Specifically, the boundedness smoothness can only guarantee
$\text{r}_{\boldsymbol{u}_{1}}(A)-\text{r}_{\boldsymbol{u}_{2}}(A)\leq m\max_{a\in A}|u_{1}(a)-u_{2}(a)|,$
the additive nature of the linear rewards setting allow us to establish a much sharper inequality:
$\text{r}_{\boldsymbol{u}_{1}}(A)-\text{r}_{\boldsymbol{u}_{2}}(A)\leq\sum_{a\in A}|u_{1}(a)-u_{2}(a)|.$
Leveraging this stronger boundedness condition, we can derive the improved regret bound (21).
Appendix D Proofs for Regret Lower Bounds
D.1  Notational Set Up and Auxiliary Results
@@
Notational Set Up. We recall the notational set-up in Section 3. Consider a non-anticipatory policy $\pi$ and an instance $I$ with reward distribution $P$. We denote $\rho_{P,\pi}$ as the joint probability distribution function on $S,A_{1},R_{1},\ldots,A_{T},R_{T}$, the concatenation of the offline dataset $S$ and the online trajectory under policy $\pi$ on instance $I$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable event $E$, we denote $\Pr_{P,\pi}(E)$ as the probability of $E$ holds under $\rho_{P,\pi}$. For a $\sigma(S,A_{1},R_{1},\ldots,A_{T},R_{T})$-measurable random variable $Y$, we denote $\mathbb{E}_{P,\pi}[Y]$ as the expectation of $Y$ under the joint probability distribution function $\rho_{P,\pi}$. In our analysis, we make use of $Y$ being $\text{Reg}_{T}(\pi,P)$ or $N_{T}(a)$ for an arm $a$. To lighten our notational burden, we abbreviate $\mathbb{E}_{P,\pi}[\text{Reg}_{T}(\pi,P)]$ as $\mathbb{E}[\text{Reg}_{T}(\pi,P)]$.
@@
Auxiliary Results To establish the regret lower bounds and the impossibility result, we need three auxiliary results, namely Claim Claim and Theorems 10, 11. Claim Claim is on the KL-divergence between Gaussian random variables:
Claim.
@@
The following, dubbed Bretagnolle–Huber inequality, is extracted from Theorem 14.2 from lattimore2020bandit:
Theorem 10.
@@
Lastly, the derivation of the chain rule in Theorem 11 largely follows from the derivation of Lemma 15.1 in lattimore2020bandit:
Theorem 11.
@@
We provide a proof to theorem 11 for completeness sake:
Proof.
Proof of Theorem 11 The proof largely follow the well-known chain rule in the multi-armed bandit literature, for example see Lemma 15.1 in lattimore2020bandit. We start by explicitly expressing the joint probability function $\rho_{P,\pi}$ on
@@
$\qquad\log\left(\frac{P^{\text{(on)}}_{a_{t}}(r_{t})}{Q^{\text{(on)}}_{a_{t}}(r_{t})}\right)\text{ d}x\text{ d}r_{1}\ldots\text{ d}r_{t}$
@@
$\qquad\times\pi_{t}(a_{t}|x,a_{1},r_{1},\ldots,a_{t-1},r_{t-1})\text{ d}x\text{ d}r_{1}\ldots\text{ d}r_{t-1}\Big\}\cdot\int_{r_{t}}P^{\text{(on)}}_{a_{t}}(r_{t})\log\left(\frac{P^{\text{(on)}}_{a_{t}}(r_{t})}{Q^{\text{(on)}}_{a_{t}}(r_{t})}\right)\text{ d}r_{t}$
@@
D.2  Proof for instance dependent regret lower bound for Multi-armed Bandit
D.2.1  Proof for Theorem 3